Please help: Solve and check answeres.
3(4-y)=5(y+2)

Answer by Nerd helper (4)

3(4-y)= 5(y+2)
Distribute first
12-3y= 5y+10
Then subtract the 5y to the left side of the equation
12-8y=10
Next subtract the 12 to the right side
-8y= -2
Finally divide the -8 on both sides
y= 1/4 or 0.25 after you simplify it